OpenStudy (anonymous):

use the intermediate value theorem to show that a solution exist for Sin(x)=x-1 in the interval [0,pi/2]

OpenStudy (zarkon):

doesn't look true

OpenStudy (anonymous):

Just had an exam I put that it isn't true but I am not sure

myininaya (myininaya):

Let \[f(x)=\sin(x)-x+1\] \[f(0)=\sin(0)-0+1=1>0\] \[f(\frac{\pi}{2})=\sin(\frac{\pi}{2})-\frac{\pi}{2}+1=2-\frac{\pi}{2}>0\] we could try a smaller interval but it is true

OpenStudy (zarkon):

it's not true...the root is at around 1.93

OpenStudy (anonymous):

right and we were not supposed to use calculators.

myininaya (myininaya):

graphing f(x)=sin(x)-x+1 i have an x-intercept on the interval [0,pi/2]

OpenStudy (anonymous):

lol I put that [0,pi/2] on the side just in case

myininaya (myininaya):

x=1.0177624

OpenStudy (anonymous):

this question was worth 20 pts maybe I'll get 5 for trying

OpenStudy (zarkon):

are you sure about that x-intercept

myininaya (myininaya):

plug it in it works

myininaya (myininaya):

i mean its an approximation so you should get approximately the same thing on both sides

OpenStudy (zarkon):

http://www.wolframalpha.com/input/?i=plot [Sin%28x%29-x%2B1]+x%3D0+to+x%3Dpi%2F2

OpenStudy (anonymous):

okay this is what I did. I used 0 and plugged it in to the equation and got -1 and than I plugged pi/2 and got stuck

OpenStudy (across):

Do you know what the IVT (intermediate value theorem) is? If so, then notice that you're trying to solve\[\sin(x)-x+1=0.\]You're given the range\[[0,\frac{\pi}{2}].\]What are the values of this function at the boundaries?\[f(0)=\sin(0)-(0)+1=1,\]\[f(\frac{\pi}{2})=\sin(\frac{\pi}{2})-(\frac{\pi}{2})+1=2-\frac{\pi}{2}.\]However,\[[1,2-\frac{\pi}{2}]\]is a range that doesn't include 0. Therefore, how can it have a solution in that interval? Perhaps I'm doing something wrong. The solution is at x ≈ 1.93456, which is greater than pi/2 ≈ 1.57079.

myininaya (myininaya):

then my calc is not working

OpenStudy (zarkon):

The IVT is inconclusive for this problem.

myininaya (myininaya):

graph sin(x)-x+1 in your calculator zarkon

myininaya (myininaya):

the ti83 plus

OpenStudy (zarkon):

it does not tell us that we do not have a root (even though we don't)

OpenStudy (zarkon):

are you in radians

myininaya (myininaya):

darn it i will never tell you

OpenStudy (zarkon):

lol

myininaya (myininaya):

ok you win you got 20/20

OpenStudy (zarkon):

I would look at the derivative of f

OpenStudy (zarkon):

f'(x)=cos(x)-1<0 on [0,pi/2]

myininaya (myininaya):

i told my calc to always be in radians why is not a faithful companion

OpenStudy (zarkon):

f(pi/2)>0 thus there can be no root

OpenStudy (zarkon):

bad calc...bad calc